x+6y=9 x=2-6y is there a solution of ordered pair or none
I think there isnt one but double checking

Answer by algebrahouse.com(1079)

x + 6y = 9
x = 2 - 6y

2 - 6y + 6y = 9 {substituted 2 - 6y, in for x, into top equation}
2 = 9 {combined like terms}
= no solution {the variable disappeared and you got a false statement}

When the variable disappears and you get a false statement, such as 0 = 3,
then thee is "no solution". This means, the two lines would not intersect.
In other words, the two equations, when graphed, would create parallel lines.

When the variable disappears and you get a true statement, such as 0 = 0,
then there are "infinite solutions". This means, the two lines intersect
in an infinite number of points. In other words, the two equations are equivalent,
and would create the same line when graphed.
For more help like this, ask the house at www.algebrahouse.com